OpenStudy (anonymous):

The bill from your plumber was $134. The cost for labor was $32 per hour. The cost for materials was $46. How many hours did the plumber work?

OpenStudy (anonymous):

32H+46=134

OpenStudy (anonymous):

32H=88

OpenStudy (anonymous):

H=2.75 (OR 2 3/4)

OpenStudy (anonymous):

h=2 and 45 minutes

OpenStudy (anonymous):

A.) define the variable for each question. B.) write an equation and c.) solve

OpenStudy (anonymous):

134-46= $88 for labor 88/32 = 2.75 hours = 2 hours 45 minutes